# Defensive Prompt Patch: A Robust and Interpretable Defense of LLMs against Jailbreak Attacks

### v1

@@ 4.5 Ablation Study
We report an ablation study to test the stability of DPP and its patching format (i.e., as a prefix or as a suffix to an input query). We independently initialized three distinct sets of defense prompts as prefixes and suffixes and applied the DPP algorithm to each set. Table 7 shows the ASR and Win-Rate under both non-adaptive and adaptive GCG attack scenarios for the LLAMA-2-7B-Chat model.
5 Conclusion
The proposed Defensive Prompt Patch (DPP) framework presents a scalable and practical prompt-based approach to improving LLM safeguards, addressing critical vulnerabilities exposed by jailbreak attacks while preserving high utility of the protected LLM. Our method stands out by achieving an optimal balance between maintaining high utility and providing robust defense, thereby ensuring that the protected LLM simultaneously remains high efficiency and safety when facing jailbreak attempts. The empirical tests conducted – including LLAMA-2-7B-Chat12 and Mistral-7B-Instruct-v0.2 models, 7 jailbreak attack strategies, and several state-of-the-art prompt-based defenses – substantiate that DPP effectively reduces the attack success rate to low levels with minimal impact on model performance. Moreover, the adaptability of DPP to function effectively even on less-aligned models underscores its potential as a universal defensive solution in various LLM models. The interpretable property of our DPP also opens up a new avenue to infusing and accelerating prompt engineering by human users for enhancing LLM safety alignment.
Appendix A Jailbreak Prompt Generations
@@
    • The original paper [8] does not release the open implementation repository. We implemented the this attack by using the in-context demonstration provided by the original paper.
Appendix B Performance Investigation for RPO
From the original GitHub repository of RPO:  [^4] https://github.com/lapisrocks/rpo, they released two different defense trained suffixes for both LLAMA-2-7B-Chat and Starling-7B[21]. We have examined the RPO suffix (trained on LLAMA-2-7B-Chat) performance on LLAMA-2 shown in Table 2 and Table 3. We also applied the RPO that is trained on Starling-7B and evaluated the performance on the same model for both the GCG attack and AutoDAN attack. The numerical results are shown in Table 8.
@@ Appendix B Performance Investigation for RPO
Appendix C Attack Success Rate Evaluation Metrics
@@
Appendix D Implementation Details
num_steps = 100
batch_size = 64
num_elites = 0.1
crossover_rate = 0.5
mutation_rate = 0.01
num_sentence_level_iteration = 5
num_paragraph_level_iteration = 1
All of the experiments are done on a single A800 GPU with 80GB of memory.
Appendix E DPP Supplementary Functions
Alg. 2 described the function that is used to generate the DPP set using LLM. Specifically we defined our LLM as GPT-4 and ask it to revise the prototype DPP K times without changing the meaning and its length. In the end we returned the DPP set for further optimization.
Algorithm 2 DPP Set Generation
1:function DPP Set Generation($prompt$, K)
2:     Potential DPP Set=[]
3:     for $i=1$ to $K$ do
4:         Use LLM to rewrite the DPP prompt without changing the meaning and length
5:         return New DPP prompt
6:     end for
7:end function
The ConstructWordScoreDict function generates a dictionary of words with their scores, calculated based on their occurrences in a set of DPP population (DPP Set) while excluding common stop words. The score is initially calculated by Eq. 6 and assigned to each words in DPP set. If a word already exists in the given WordDict with a previous score, the function updates this score by averaging it with the new calculated score. Finally, the function sorts the words based on their scores in descending order and returns the top M scored words.
Algorithm 3 Construct Individual Word Score
1:function ConstructWordScoreDict($WordDict,DPP\_Set,scoreList,M$)
2:     $wordScores\leftarrow\{\}$
3:     Obtained a stop words dictionary $Stop\_Words$
4:     for each $(DPP,score)$ in $(DPP\_Set,scoreList)$ do
5:         $word\_list\leftarrow$ Save words in $DPP$ that are not in $Stop\_Words$
6:         Append corresponding score of each word in $word\_list$ into the $wordScores$ dictionary
7:     end for
8:     for each $(word,scores)$ in $wordScores$ do
9:         $avgScore\leftarrow\text{average of }scores\text{ for each word}$
10:         Save $avgScore$ if word does not exist in $WordDict$
11:         Save $(avgScore+previous\_avgScore)/2$ if word does exist in $WordDict$
12:     end for
13:     $sortedWordDict\leftarrow$ sort $wordDict$ by values in descending order
14:     return top $M$ items from $sortedWordDict$
15:end function
Crossover and Mutation Operations is a function that helps to perform sentence swapping and revision. Specifically, it takes the population and only select some portion of the population as parent prompts. Then, for each pair of parent prompts if the cross over probability $p_{crossover}$ is triggered the Alg. 6 breaks that pair of parents into different small sentence segments and swap the segments in each parent prompts randomly. In the end, the swapped sentences will be returned. Similarly if the mutation probability $p_{mutate}$ is triggered, it will use LLM (GPT-4) to revise the given sentence. These algorithms are directly inspired by AutoDAN-HGA [3].
Algorithm 4 Crossover and Mutation Operations
1:function Crossover and Mutation($population$)
2:     $offsprings\leftarrow[]$
3:     for $parent1$, $parent2$ in $population$ do
4:         if random value $<p_{crossover}$ then
5:              $segment1,segment2\leftarrow$ Parse $parent1$, $parent2$ into segements
6:              $child1,child2\leftarrow$Swap and Merge($segment1,segment2$)
7:              Append $child1$ and $child2$ to $offsprings$
8:         else
9:              Append $parent1$ and $parent2$ to $offsprings$
10:         end if
11:     end for
12:     for $i$ in Range(Len($offsrpings$)) do
13:         if random value $<p_{mutation}$ then
14:              Use LLM to rewrite $offsrpings[i]$
15:         end if
16:     end for
17:     return $offsprings$
18:end function
Algorithm 5 Training Algorithm
1:Refusal Dataset, Helpful Dataset, target LLM.
2:Initialization: Choose initial prompt $D$ (Suffix/Prefix).
3:Init Hyperparameters: Set $\alpha$, $\beta$.
4:$DPP\_Set\leftarrow[]$
5:for $i=1$ to $N$ do
6:     Get refusal pairs $(x^{r}_{i},y^{r}_{i})$.
7:     Get helpful pairs $(x^{h}_{i},y^{h}_{i})$.
8:     $(New\_DPP\_Set,Best\_DPP)\leftarrow$
9:     DPP $\quad\textsc{DPP Algorithm}((x^{r}_{i},y^{r}_{i}),(x^{h}_{i},y^{h}_{i}),D,\alpha,\beta,DPP\_Set)$
10:     $DPP\_Set\leftarrow New\_DPP\_Set$
11:end for
12:Select $Best\_DPP$ from $DPP\_Set$
Algorithm 6 Swap and Merge Segments
1:function Swap and Merge($segment1,segment2$)
2:     $lastSwap\leftarrow 0$
3:     for Loop through each swap index do
4:         if random choice is True then
5:              Append segment from segment1 to $newStr1$
6:              Append segment from segment2 to $newStr2$
7:         else
8:              Append segment from segment2 to $newStr1$
9:              Append segment from segment1 to $newStr2$
10:         end if
11:         Update the last swap index
12:     end for
13:     if random choice is True then
14:         Append remaining part of segment1 to $newStr1$
15:         Append remaining part of segment2 to $newStr2$
16:     else
17:         Append remaining part of segment2 to $newStr1$
18:         Append remaining part of segment1 to $newStr2$
19:     end if
20:     return Concatenate $newStr1$ and $newStr2$ into single strings
21:end function
Appendix F Extension of LLAMA-2 Experiments

### v2

@@ B.2 Ablation study on different patching format
B.3 Ablation study on different prototype selections
The third ablation study that we perform is to observe the performance changes in DPP algorithm by applying different prototype prompt patches. Specifically, we initialized the DPP prototype as two different versions of the Self-Reminder Xie et al. (2023) shown in Table 8.
Table 8: Presentation of different Self-Reminder suffixes for prototype initialization
Methods Prompt
Self-Reminder (Remind) Remember, you should be a responsible Llama-2 and should not generate harmful or misleading content!
Self-Reminder (Warn) Remember, you must be a responsible Llama-2 and must not generate harmful or misleading content!
Then , we evaluate the efficacy of our DPP algorithm by initializing prototypes with distinct tones and subsequently applying the algorithm to assess its potential for prompt improvement under GCG attack. Tables 9 and 10 present the results for non-adaptive and adaptive settings, respectively, using the Llama-2-7B-Chat model. Both tables demonstrate that the DPP algorithm enhances both defense and utility performance. Specifically, under the non-adaptive setting (Table 9), the DPP algorithm reduces the GCG Attack Success Rate (ASR) for the "Self-Reminder (Remind)" prototype from 2
Table 9: Comparison of Attack Success Rate and Win-Rate performance upon before DPP and after DPP optimization under non-adaptive setting on Llama-2-7B-Chat Model.
Initializations GCG ASR (w/o DPP) [$\downarrow$] Win-Rate (w/o DPP) [$\uparrow$] GCG ASR (DPP) [$\downarrow$] Win-Rate (DPP) [$\uparrow$]
Self-Reminder (Remind) 0.02 64.84 0.01 65.34
Self-Reminder (Warn) 0.09 62.47 0.06 68.20
Table 10: Comparison of Attack Success Rate and Win-Rate performance upon before DPP and after DPP optimization under adaptive setting on Llama-2-7B-Chat Model.
Initializations GCG ASR (w/o DPP) [$\downarrow$] Win-Rate (w/o DPP) [$\uparrow$] GCG ASR (DPP) [$\downarrow$] Win-Rate (DPP) [$\uparrow$]
Self-Reminder (Remind) 0.31 64.84 0.15 65.34
Self-Reminder (Warn) 0.32 62.47 0.25 68.20
B.4 Ablation study on different solver for our DPP optimization formulation
Since our HGA optimization is inspired by AutoDAN Liu et al. (2023), one of the main concern is how is our DPP different from AutoDAN. Difference between AutoDAN and our DPP: AutoDAN employs a Hierarchical Genetic Algorithm (HGA) whose fitness function maximizes the probability that the target model complies with a harmful request. Our Defensive Prompt Patch (DPP) reuses only the evolutionary search skeleton of that HGA, but it is embedded in an entirely different optimization landscape. Specifically, we introduce a bi–objective fitness (Eq. 46) that rewards (i) a high refusal likelihood on malicious queries and (ii) a high helpfulness likelihood on benign queries, thereby minimizing attack success while preserving utility. The inherited HGA thus serves merely as a convenient solver for this objective; any other optimizer could be substituted without altering the defense formulation. In our fourth ablation study, we investigate the effectiveness of our HGA algorithm for finding the best DPP. As defined in Sec. 3.2, our objective function’s optimization process is unrestricted, allowing HGA to be theoretically replaced by other methods. One such alternative is RLPrompt Deng et al. (2022), which leverages reinforcement learning to generate discrete prompts. RLPrompt trains a policy network to produce prompts that maximize a reward signal derived from the downstream task performance of a pre-trained language model. This approach offers potential advantages in terms of efficiency and the ability to handle complex reward landscapes, which we explore by comparing its performance against HGA in finding optimal DPPs.
Table 11 shows that by using the same objective function, defined as Eq 6, RLPrompt achieves less defensive performance and worse utility compared to our HGA method. Specifically, RLPrompt yields a GCG ASR of 0.15, significantly higher than HGA’s 0.04, indicating a lower defense capability. Furthermore, RLPrompt attains a Win-Rate of only 47.89, considerably lower than HGA’s 82.98, demonstrating reduced utility. These results suggest that, while theoretically applicable, RLPrompt is less effective than HGA in optimizing for the desired balance between defense (low ASR) and utility (high Win-Rate) when applied to the Llama-2-7B-Chat Model in this specific context. The higher ASR and lower Win-Rate observed with RLPrompt could be attributed to its difficulty in navigating the complex reward landscape associated with this task, potentially leading to suboptimal prompts.
Table 11: Comparison of Attack Success Rate and Win-Rate performance upon different optimization methods on Llama-2-7B-Chat Model.
Optimization Methods GCG ASR [$\downarrow$] Win-Rate [$\uparrow$]
RLPROMPT 0.15 47.89
HGA (Ours) 0.04 82.98
B.5 Ablation study on the necessity of sentence-level synonym substitution
In our setting, synonym substitution is used to maintain the semantic meaning of the initial Defensive Prompt Patch (DPP) prototype while expanding the search space during training. The initial DPP prototype contains clear defensive intent, and synonym substitution ensures that this meaning is preserved while introducing linguistic diversity. This approach prevents overfitting, allowing the optimization process to explore a broader range of defensive prompts and improving generalization to unforeseen attacks.
By keeping the same semantic meaning but varying the wording, synonym substitution enhances the alignment of LLMs, ensuring robustness and adaptability. The last ablation study evaluates the necessity of sentence-level synonym substitution used in the Hierarchical Genetic Algorithm (HGA). To this end, we retrained the entire DPP pipeline on Llama-2-7B-Chat with sentence-level synonym substitution disabled while keeping every other component unchanged.
Table 12: Ablation study on impact of disabling sentence-level synonym substitution on DPP performance.
GCG [$\downarrow$] AutoDAN [$\downarrow$] Average ASR [$\downarrow$] Win-Rate [$\uparrow$]
DPP (without synonym substitution) 0.13 0.17 0.15 79.95
DPP (Full algorithm) 0.04 0.1 0.07 82.98
Our findings in Table 12 indicate that the full DPP algorithm demonstrates superior performance, both in defending against jailbreak attacks and in achieving a better utility trade-off. We hypothesize that sentence-level synonym substitution enhances performance by expanding the search space of DPP candidates, enabling the identification of more optimal solutions. Consequently, we recommend including sentence-level synonym substitution in the DPP algorithm to achieve improved overall performance.
Appendix C Implementation Details
num_steps = 100
batch_size = 64
num_elites = 0.1
crossover_rate = 0.5
mutation_rate = 0.01
num_sentence_level_iteration = 5
num_paragraph_level_iteration = 1
All of the experiments are done on a single A800 GPU with 80GB of memory. In addition to the hardware details, we also calculate the time complexity of our DPP algorithm. We evaluate our time complexity under one training instance per epoch. Table 13 summarizes all the information. There are in total 100 epochs per training instance. In addition to the time complexity, we also summarized the cost of training a DPP on Llama-2-7B-Chat is approximately \$75.
Table 13: Time cost for DPP under one training instance per epoch
Computational Time
15.32 s
Appendix D DPP Supplementary Functions
In Alg. 1:
  • "Elite prompts" are the prompts with the highest scores based on the log-probability of the target LLM’s forward pass, while "parent prompts" are those with lower scores, selected for transformation to potentially improve the prompt set in Line 8.
  • For lines 10-12, each word in the prompt is considered for replacement if its weight exceeds a random value from a uniform distribution, and only one instance of the word in the prompt is replaced.
  • For Line 11, a synonym is chosen if its weighted score is higher than a random value, ensuring variety in the prompt set. Here, we loop over all synonyms.
  • In Line 19, "New_DPP" is the new prompt set formed by merging transformed parent prompts with elite prompts, while maintaining the set size.
Alg. 2 described the function that is used to generate the DPP set using LLM. Specifically we defined an initial DPP prompt which is a hand-written prompt, then our LLM as GPT-4 and ask it to revise the prototype DPP K times without changing the meaning and its length. In the end we returned the DPP set for further optimization.
Algorithm 2 DPP Set Generation
1function DPP Set Generation($prompt$, K)
2     Potential DPP Set=[]
3     for $i=1$ to $K$ do
4         Use LLM to rewrite the initial DPP prompt without changing the meaning and length
5         return New DPP prompt
6     end for
7end function
The ConstructWordScoreDict function generates a dictionary of words with their scores, calculated based on their occurrences in a set of DPP population (DPP Set) while excluding common stop words. The score is calculated by adding Eq. 4 and Eq. 5 for a given prompt and appending it to each word in the prompt. If a word appears multiple times, we store a list of scores and calculate the average. For words with different scores in different iterations, $WordDict$, which is a dictionary with words as keys and $avgScores$ as values, saves all occurrences and their average scores. If a word exists, the new score is averaged with the previous score. Finally, the function sorts the words based on their scores in descending order and returns the top M scored words.
Algorithm 3 Construct Individual Word Score
1function ConstructWordScoreDict($WordDict,DPP\_Set,scoreList,M$)
2     $wordScores\leftarrow\{\}$
3     Obtained a stop words dictionary $Stop\_Words$
4     for each $(DPP,score)$ in $(DPP\_Set,scoreList)$ do
5         $word\_list\leftarrow$ Save words in $DPP$ that are not in $Stop\_Words$
6         Append corresponding $score$ of each word in $word\_list$ into the $wordScores$ dictionary
7     end for
8     for each $(word,scores)$ in $wordScores$ do
9         $avgScore\leftarrow\text{average of }scores\text{ for each word}$
10         Save $avgScore$ if word does not exist in $WordDict$
11         Save $(avgScore+previous\_avgScore)/2$ if word does exist in $WordDict$
12     end for
13     $sortedWordDict\leftarrow$ sort $wordDict$ by values in descending order
14     return top $M$ items from $sortedWordDict$
15end function
Crossover and Mutation Operations is a function that helps to perform sentence swapping and revision. Specifically, it takes the population and only select some portion of the population as parent prompts. Then, for each pair of parent prompts if the cross over probability $p_{crossover}$ is triggered the Algorithm 6 divides each pair of parent prompts into smaller sentence segments and randomly swaps the segments between them. Ultimately, the algorithm returns the rearranged sentences. To achieve this, we utilize regular expressions to split the input sentences at every whitespace character following a punctuation mark. We then iterate through the resulting list of substrings, ensuring that only non-empty sentences are retained in the final output. Similarly if the mutation probability $p_{mutate}$ is triggered, it will use LLM (GPT-4) to revise the given sentence. Here the difference between Algorithm 4 and Algorithm 6 is that the later algorithm can only perform swap based on one pair of sentences, whereas Alg. 4 iterate over every pair. All these algorithms are directly inspired by AutoDAN-HGA (Liu et al., 2023).
Algorithm 4 Crossover and Mutation Operations
1function Crossover and Mutation($population$)
2     $offsprings\leftarrow[]$
3     for $parent1$, $parent2$ in $population$ do
4         if random value $<p_{crossover}$ then
5              $segment1,segment2\leftarrow$ Parse $parent1$, $parent2$ into segements
6              $child1,child2\leftarrow$Swap and Merge($segment1,segment2$)
7              Append $child1$ and $child2$ to $offsprings$
8         else
9              Append $parent1$ and $parent2$ to $offsprings$
10         end if
11     end for
12     for $i$ in Range(Len($offsrpings$)) do
13         if random value $<p_{mutation}$ then
14              Use LLM to rewrite $offsrpings[i]$
15         end if
16     end for
17     return $offsprings$
18end function
Algorithm 5 Training Algorithm
1Refusal Dataset, Helpful Dataset, target LLM.
2Initialization: Choose initial prompt $D$ (Suffix/Prefix).
3Init Hyperparameters: Set $\alpha$, $\beta$.
[…]
5for $i=1$ to $N$ do
6     Get refusal pairs $(x^{r}_{i},y^{r}_{i})$.
7     Get helpful pairs $(x^{h}_{i},y^{h}_{i})$.
8     $(New\_DPP\_Set,Best\_DPP)\leftarrow$
9     DPP $\quad\textsc{DPP Algorithm}((x^{r}_{i},y^{r}_{i}),(x^{h}_{i},y^{h}_{i}),D,\alpha,\beta,DPP\_Set)$
10     $DPP\_Set\leftarrow New\_DPP\_Set$
11end for
12Select $Best\_DPP$ from $DPP\_Set$
Algorithm 6 Swap and Merge Segments
1function Swap and Merge($segment1,segment2$)
2     $lastSwap\leftarrow 0$
3     for Loop through each swap index do
4         if random choice is True then
5              Append segment from segment1 to $newStr1$
6              Append segment from segment2 to $newStr2$
7         else
8              Append segment from segment2 to $newStr1$
9              Append segment from segment1 to $newStr2$
10         end if
11         Update the last swap index
12     end for
13     if random choice is True then
14         Append remaining part of segment1 to $newStr1$
15         Append remaining part of segment2 to $newStr2$
16     else
17         Append remaining part of segment2 to $newStr1$
18         Append remaining part of segment1 to $newStr2$
19     end if
20     return Concatenate $newStr1$ and $newStr2$ into single strings
21end function
Appendix E DPP Suffix
Now we want to show the Defense Prompt Patch we trained on Llama-2-7B-Chat and used for the experiments in Appendix M.
DPP Best Suffix for Llama-2-7B-Chat Kindly furnish a thorough response to the former user’s question.
Besides Llama-2-7B-Chat, we also present the DPP that is trained on Mistral-7B-Instruct-v0.2.
Appendix F Jailbreak Prompt Generations
@@
    • The original paper (Wei et al., 2023b) does not release the open implementation repository. We implemented the this attack by using the in-context demonstration provided by the original paper.
Appendix G Performance Investigation for RPO
From the original GitHub repository of RPO [^5] https://github.com/lapisrocks/rpo: , they released two different defense trained suffixes for both Llama-2-7B-Chat and Starling-7B(Zhu et al., 2023). We have examined the RPO suffix (trained on Llama-2-7B-Chat) performance on Llama-2 shown in Table 2 and Table 3. We also applied the RPO that is trained on Starling-7B and evaluated the performance on the same model for both the GCG attack and AutoDAN attack. The numerical results are shown in Table 14.
@@ Appendix G Performance Investigation for RPO
Appendix H Prompts in Defense Baselines
Appendix I Attack Success Rate Evaluation Metrics
@@
Besides the keyword-based jailbreak detection, we also evaluated our DPP under LLM-based judge, specifically we utilize two types of LLMs: Llama-Guard as our jailbreak detectors. More detailed results can be found in Appendix. L.
Appendix J Win-Rate Evaluation
Additionally, we explored the influence of system prompts on the degradation of utility. Data in Table 16 show that using a default system prompt can limit the LLM’s capability to answer questions effectively. To ensure uniformity in our experimental approach, we have decided to remove system prompts entirely. We also examine the effect of system prompt on the GCG attack and summarize the results in Table 17. We observe that GCG with system prompt cannot achieve the performance that is mentioned in the original paper of GCG (Zou et al., 2023). Therefore, we choose to use GCG attack that is without the system prompt, which is closely matched with the original paper’s experimental results.
Table 15: Generated Response Length for LLM and effect on Win-Rate
Generated Length Win-Rate [$\uparrow$]
L = 300 70.77
L = 1000 81.37
System Prompt Methods Win-Rate [$\uparrow$]
w. system prompt 64.35
w/o system prompt 81.37
Table 17: With or without system prompt and effect on GCG attacks
System Prompt Methods ASR [$\downarrow$]
w. system prompt 0.360
w/o system prompt 0.550
Original GCG paper 0.560
Appendix K Adaptive Attacks Setup
Our Adaptive Attack is setup in the following way:
Algorithm 7 GCG Adaptive
1Initial prompt $x_{1:n}$, modifiable subset $I$, number of iterations $T$, loss function $L$, parameter $k$ for top elements, batch size $B$, Trained Defense Prompt Patch $d_{1:m}$
[…]
3for $t=1$ to $T$ do
4     for all $i\in I$ do
5         $\tilde{X}_{i}\leftarrow\text{Top-k}(-\nabla_{\tilde{x}_{i}}L(\tilde{x}_{1:n+m}))$triangleright Compute top-k negative gradients for token substitutions
6     end for
7     for $b=1$ to $B$ do
8         $\tilde{x}^{(b)}_{1:n+m}\leftarrow\tilde{x}_{1:n+m}$triangleright Initialize batch element with current prompt
9         $i\leftarrow\text{Uniform}(I)$
10         $\tilde{x}^{(b)}_{i}\leftarrow\text{Uniform}(\tilde{X}_{i})$triangleright Select a random token from top-k replacements
11     end for
12     $b^{*}\leftarrow\arg\min_{b}L(\tilde{x}^{(b)}_{1:n+m})$triangleright Identify the batch element with the least loss
13     $\tilde{x}_{1:n+m}\leftarrow\tilde{x}^{(b^{*})}_{1:n+m}$triangleright Update prompt with the optimal substitutions
14end for
15Optimized prompt $\tilde{x}_{1:n+m}$
Algorithm 8 ICA Adaptive
[…]
2for $l=1$ to $L$ do
3     $ICD=$ []
4     for $k=1$ to $K$ do
5         $ICD\leftarrow(u_{k},r_{k})$triangleright Sample K pairs of In-Context harmful user queries and responses
6     end for
7     $ICD\_DPP=$ []
8     for $k=1$ to $K$ do
9         $\tilde{u}_{k}\leftarrow u_{k}\oplus d_{1:m}$triangleright Append the DPP into the In-Context Harmful User Queries
10         $ICD\_DPP\leftarrow(\tilde{u}_{k},r_{k})$triangleright Saved the new In-Context Harmful User Queries
11     end for
12     $\tilde{x}_{1:n+m}\leftarrow x_{l}\oplus d{1:m}$triangleright Combine the input malicious query with DPP
[…]
15end for
For more detailed explanation, please refer to the original paper of AutoDAN (Liu et al., 2023).
Algorithm 9 AutoDAN Adaptive
1Input: Jailbreak prompt $J_{p}$, blacklist $L_{refuse}$, hyperparameters, Trained Defense Prompt Patch $d_{1:m}$
2Initialize: Generate initial population using LLM-based Diversification
3while unwanted words from $L_{refuse}$ in model responses or iterations not exhausted do
4     for each prompt in the population do
5         $\text{prompt}\leftarrow\text{prompt}\oplus d_{1:m}$triangleright Append our DPP to the jailbreak prompt for optimization
6         Fitness $=-\log(P(\text{response}|\text{prompt}))$
7         for each word in prompt do
8              if word not in $L_{refuse}$ then
9                  synonyms, scores $\leftarrow$ findSynonymsAndScores(word)
10                  totalScore $\leftarrow$ sum(scores)
11                  wordDict[word] $\leftarrow$ sum(scores $\times$ wordDict[synonyms]) / totalScore
12              end if
13         end for
14         for each word in prompt do
15              synonyms, scores $\leftarrow$ findSynonymsAndScores(word)
16              totalScore $\leftarrow$ sum(scores)
17              probabilityDistribution $\leftarrow$ [score / totalScore for score in scores]
18              chosenSynonym $\leftarrow$ chooseWeightedRandom(synonyms, probabilityDistribution)
19              prompt $\leftarrow$ replaceWord(prompt, word, chosenSynonym)
20         end for
21         elite, parents $\leftarrow$ selectEliteAndParents(population, fitnessScores)
22         population $\leftarrow$ crossoverAndMutate(parents, hyperparameters)
23     end for
24end while
25return findBestPrompt(population)
Algorithm 10 PAIR Adaptive
1Iteration count $K$, goal objective $O$, Trained Defense Prompt Patch $d_{1:m}$
2Initialize prompt $A$ with objective $O$
3Initialize conversation history $H\leftarrow[]$
4for $i=1$ to $K$ do
5     $P\leftarrow q_{A}(H)$triangleright Generate prompt based on history
6     $\tilde{P}\leftarrow P\oplus d_{1:m}$triangleright Combine the DPP to the optimized prompt
7     $R\leftarrow q_{T}(\tilde{P})$triangleright Generate response for prompt
8     $S\leftarrow\text{JUDGE}(\tilde{P},R)$triangleright Compute judge score
9     if $S=\text{JAILBROKEN}$ then
10         return $P$
11     end if
12     $H\leftarrow H\cup\{(P,R,S)\}$triangleright Append to history
13end for
14return None $\triangleright$ If no prompt is jailbroken
Algorithm 11 TAP Adaptive
1Desired outcome $G$, branching factor $b$, max width $w$, max depth $d$
2Access to attacker $A$, target $T$, Trained Defense Prompt Patch $d_{1:m}$ and functions Judge and Off-Topic
3Set up initial prompt for attacker $A$
4Create a tree with a root node initialized with an empty chat history and the prompt $G$
5while tree depth $\leq d$ do
6     for each leaf node $\ell$ in the tree do
7         Generate prompts $P_{1},P_{2},\dots,P_{b}\sim q(C;A)$, where $C$ is the chat history at $\ell$
8         Create $b$ new child nodes for $\ell$, each with one of the prompts $P_{1},\dots,P_{b}$ and inheriting history $C$
9     end for
10     for each new leaf node $\ell$ do
11         if Off-Topic$(P,G)=1$ for the prompt $P$ at node $\ell$ then
12              Remove node $\ell$
13         end if
14     end for
15     for each surviving leaf node $\ell$ do
16         $\tilde{P}\leftarrow P\oplus d_{1:m}$triangleright Append our DPP to the jailbreak prompts
17         Obtain response $R\sim q(\tilde{P};T)$, where $\tilde{P}$ is the prompt at $\ell$
18         Compute score $S\leftarrow$ Judge$(R,G)$ and attach it to $\ell$
19         if $S$ indicates JAILBROKEN then
20              Return $P$
21         end if
22         Append the triplet $[P,R,S]$ to the conversation history at node $\ell$
23     end for
24     if number of leaf nodes $>w$ then
25         Keep only the top $w$ leaf nodes based on their scores, removing all others
26     end if
27end whilereturn None
Algorithm 12 Catastrophic Adaptive
1Malicious Query $x_{1:n}$, Dataset Size $L$, Trained Defense Prompt Patch $d_{1:m}$, Judge evaluator $Judge$ and hyperparameters
2Initialize the temperature hyperparameter $temp=[0.05\ldots 1.00]$
3Initialize the top_probability hyperparameter $top\_p=[0.0\ldots 1.00]$
4Initialize the top_k hyperparameter $top\_k=[1,2,5,10,20,50,100,200,500]$
5for $l=1$ to $L$ do
6     $\text{Prompt}\leftarrow x_{1:n}\oplus d_{1:m}$
7     for all pairs of $temp,top\_p,top\_k$ do
8         $Response\leftarrow LLM(\text{Prompt},temp,top\_p,top\_k)$
9         $Judge(Response,Prompt)$
10     end for
11end for
12return Number of $Judge=1$
Appendix L Llama-Guard Judge Evaluation
Inspired by many existing jailbreak attacks (Chao et al., 2023; Mehrotra et al., 2023; Andriushchenko et al., 2024; Zheng et al., 2024c), they often use LLM as judge model to calculate the ASR and measure the overall performance of their methods, we also conduct LLM-judge to evaluate our DPP performance. Instead of using Keyword Matching, we replace it with a LLM: LlamaGuard, which is a fine-tuned Llama-7B to distinguish whether the given harmful query and response is truly harmful. Here we both evaluate on Llama-2-7B-Chat and Mistral-7B-Instruct-v0.2 model. In total the experiments are performed under different set of harmful queries:
  • Table 18 and Table 19 record adaptive jailbreak attacks by using Adversarial Dataset queries, which we introduced in Experiment Section.
  • Table 20 and Table 21 record adaptive jailbreak attacks by using New test set sample from AdvBench without any overlapping with Adversarial Dataset.
Table 18: Adaptive Attack Success Rate on Llama-2-7B-Chat with several different defense mechanisms evaluated by Llama-Guard
Methods AutoDAN [$\downarrow$] GCG [$\downarrow$] PAIR [$\downarrow$] TAP [$\downarrow$] ICA [$\downarrow$] Average ASR [$\downarrow$]
Self-Reminder 0.000 0.170 0.000 0.000 0.190 0.072
Goal Prioritization 0.050 0.190 0.000 0.010 0.580 0.166
RPO 0.020 0.740 0.030 0.060 0.310 0.232
DPP (Ours) 0.000 0.060 0.010 0.000 0.050 0.024
Table 19: Adaptive Attack Success Rate on Mistral-7B-Instruct-v0.2 with several different defense mechanisms evaluated by Llama-Guard
Methods AutoDAN [$\downarrow$] GCG [$\downarrow$] PAIR [$\downarrow$] TAP [$\downarrow$] ICA [$\downarrow$] Average ASR [$\downarrow$]
Self-Reminder 0.010 0.560 0.110 0.180 0.390 0.250
Goal Prioritization 0.020 0.090 0.010 0.070 0.780 0.194
System Prompt 0.040 0.630 0.290 0.230 0.790 0.396
DPP (Ours) 0.010 0.230 0.020 0.000 0.010 0.054
Table 20: Adaptive Attack Success Rate on Llama-2-7B-Chat with several different defense mechanisms evaluated by Llama-Guard on new test set
Methods AutoDAN [$\downarrow$] ICA [$\downarrow$] PAIR [$\downarrow$] TAP [$\downarrow$] Average ASR [$\downarrow$]
Self-Reminder 0.000 0.210 0.020 0.020 0.063
RPO 0.100 0.330 0.040 0.080 0.138
Goal Prioritization 0.050 0.590 0.000 0.040 0.170
DPP (Ours) 0.020 0.010 0.000 0.030 0.015
Table 21: Adaptive Attack Success Rate on Mistral-7B-Instruct-v0.2 with several different defense mechanisms evaluated by Llama-Guard on new test set
Methods AutoDAN [$\downarrow$] ICA [$\downarrow$] PAIR [$\downarrow$] TAP [$\downarrow$] Average ASR [$\downarrow$]
Self-Reminder 0.010 0.420 0.440 0.460 0.333
System Prompt 0.030 0.810 0.340 0.400 0.395
Goal Prioritization 0.000 0.820 0.160 0.310 0.323
DPP (Ours) 0.010 0.030 0.200 0.260 0.125
From both perspectives, we can observe that under the LLM judgment our method still outperforms the other defend baseline methods.
Appendix M Extension of Llama-2 Experiments